# Effective Note-taking and its Impact on Learning Undergraduate Introductory Physics Courses

### v1

@@ III.2 Grading Policy
FSU follows the FN grade policy, which means if students do not show up in the first week of the class, they will automatically drop out of the class after the instructor notifies the registration office. Students who dropped out from getting FN grades are not counted in the university census. UNCP also follows a similar policy. There was no mid-semester withdrawal in both courses in the FSU and UNCP, although some students did not complete many assignments, which impacted their grades. As students did not voluntarily withdraw and universities do not have a policy of dropping students because of such issues, these students’ scores are included in the analysis.
III.3 Study Design
@@
Keeping the other factors the same, we implemented teaching pedagogy differing with the preparation and use of note-taking in two sets of sections. In each university, we taught effective note-taking methods, asked students to prepare handwritten notes, and let them use the notes during classwork, timed quizzes, and exams. Study shows that hand-written notes increase comprehension, create a personal connection to the learning materials, improve memory and recall of the facts, and activate brain memory areas more robustly than preparing notes using digital devices, showing the pen mightier than the Keyboard, although taking notes on electronic devices is faster, and the note taker can write to a larger extent of word-for-word what the teacher is saying in a class [51]. Students using electronic devices such as iPads, tablets, and laptops to take notes are usually prone to copying, pasting, and collecting information without a deep thought about what they are writing. In contrast, hand-written note-takers usually mentally process the information further before they write it down; while they write, they get their fingers’ extra muscle memory. Therefore, this research is designed as a comparative study of the effectiveness of handwritten notes in enhancing attention and effectiveness in physics learning. In order to maintain consistency, the quizzes and exams administered for both groups were problem-solving based, which demand critical thinking, creativity, and decision-making. Keeping all other conditions and class structure unaltered, the methodology was split into two main categories.
@@
      This policy was implemented in College Physics I (PHY 1500) at UNCP and General Physics I (PHY 115) at FSU, where the note-taking strategy was not discussed, nor were students allowed to use their notes in the quizzes and exams. However, a list of all required mathematical expressions covered in the chapters was provided during timed assignments and exams.
@@
    Handwritten note-taking is encouraged and permitted for use during exams:
      The note-taking policy was implemented in College Physics II (PHY 1510) at UNCP and General Physics II (PHY 116) at FSU. At the beginning of the semester, a few samples of the note-taking skeleton were provided, and instructors consistently observed the notes students were taking during the class for their effectiveness and guided them as per need. Only the handwritten notes prepared by students were permitted during quizzes and exams. The notes students used in the timed quizzes and exams were monitored.
In the section where the note-taking strategy was implemented, a few samples of skeleton notes of each kind discussed above, along with sample notes, were also provided in the very first week of the semester. In addition, students received timely feedback on their notes, and those who struggled with note-takings and other academic things were encouraged to use office hours for additional support. All instructors have at least eight regularly scheduled office hours per week to help students in various aspects, including course contents, their academic and career plans, and other academic support. Instructors asked students to make mind maps connecting different concepts in all the chapters covered as note-taking homework, which instructors found very helpful for students’ overall understanding of the course.
Table 1: FSU’s University catalog to assign the final letter grade for earned grade points on a scale of 0–100 and a GPA out of 4.0.
Points 90 – 100 80 – 89.9 70 – 79.9 60 – 69.9 $<$ 60
Grade A B C D F
GPA 4.0 3.0 2.0 1.0 0.0
Table 2: The UNCP’s University catalog to assign the final letter grade for earned grade points on a scale of 0–100 and a GPA out of 4.0.
Points 93–100 90–92.99 87-89.99 83-86.99 80-82.99 78-79.99 74-77.99 70-73.99 68-69.99 64-67.99 60-63.99 $<$ 60
Grade A A- B+ B B- C+ C C- D+ D D- F
GPA 4.0 3.7 3.3 3.0 2.7 2.3 2.0 1.7 1.3 1.0 0.7 0.0
@@ IV Results and discussion
The grade that each student receives is the grade that students earn based on their academic performance, not measured against the performance of other students. The final letter grades are assigned according to the respective University’s catalogs, which are tabulated in Table 1 and 2. At FSU, using the traditional five-grade system, a letter grade of ”A” (GPA of 4.0 out of 4.0) is assigned if a student receives 90% of possible credits or more than that. Letter grades ”B” (GPA of 3.0 out of 4.0), ”C” (GPA of 2.0 out of 4.0), and ”D” (GPA of 1.0 out of 4.0) are assigned if a student’s grade point lies in the range of 80 to 89.9%, 70 to 79.9%, and 70 to 79.9%, respectively. A student fails the course if the student receives less than 60% of the total possible points. In the UNCP, the letter grade assigning catalog is not a traditional five-grade, but it follows a sophisticated letter grade and plus-minus system (see Table 2).
@@
Figure 2: Bar chart showing letter grade percentage in sections with no note-taking implemented. If only the exam were considered, students’ grades and the final grade, incorporating all evaluating components included, would be presented in different colors. There were a total of 32 students: 20 from FSU and 12 from UNCP
Figure 3: Bar chart showing letter grade percentage in sections where students took notes and were allowed to use their notes during times assignments and exams. Students’ grades, if only the exam were considered and the final grade, incorporating all evaluating components included, would be presented in different colors. There were a total of 35 students: 23 from FSU and 12 from UNCP
@@
Table 3: Statistical measures of student performance in two different approaches to teaching introductory physics. The highest, lowest, and average scores were presented on the scale of 0–100.
Note-taking policy ($\rightarrow$) $Withoutnote-taking$ $Withnote-taking$
[…]
Highest 96.596.596.5 96.7 97.597.597.5 99.1
Lowest 44.544.544.5 51.5 39.439.439.4 41.7
Average 70.070.070.0 79.9 76.676.676.6 83.7
Standard deviation 16.616.616.6 13.4 14.514.514.5 12.8
At UNCP, each letter grade comes with three variations of a letter, followed by plus, letter only, and letter followed by a minus sign, except for the letter grade ”A,” which has only two variations of A and A-, and letter grade F (the failing grade) with no variation. Note that the plus and minus designation window for letter grades B and C are not the same, while the distribution range is similar for C and D. As the collective grade point range associated with a letter grade and its plus and minus at UNCP is the same to that of the letter grade at FSU, for a consistent analysis, we counted each letter along with its plus and minus at UNCP as the corresponding letter grade’s frequency such as sum of A- and A’s frequencies at UNCP is equivalent to A’s frequency at FSU, sum of B-, B+, and B’s frequencies at UNCP is equivalent to B’s frequency at FSU, and so on. Figures 2 and 3 show the percentage of letter grades that students received in sections with no note-taking and sections with note-taking were implemented. he sample size for these two groups in the study is almost the same, one being 32 and the other being 35. To better understand the effect of effective note-taking, the grades students receive when only exam scores are considered in assigning grades are presented separately. This analysis shows that homework, online quizzes, class quizzes, and other assignments helped improve students’ grades. Almost one-third of the students would fail in the sections with no note-taking implemented if only exams were considered while assigning the letter grade. Yet, nearly one-tenth of the students failed the course, even with the inclusion of all grading components. The data show that when students were allowed to take notes, the percentage of students receiving better letter grade (by better letter grade, we mean ”A”) did not change significantly. The percentage of students getting the letter grade ”A” is almost the same (please see and compare the pair of bars for letter grade ”A” of Fig. 2 with the same of Fig. 3) . The slight percentage difference is primarily due to the fact that the sample sizes are close, but not identical. While the same is not true for the percentage of students getting the letter grade ”B” and lower. With note-taking, exam-only grade follows the bell curve, peaking at B and having a tail towards a lower grade. The letter grade significantly improved when students were allowed to prepare and use the notes they had prepared. Furthermore, the percentage of students who failed the course reduced to 2.9% only. Our observation is that note-taking helped average-performing (students getting B and C letter grades), less motivated, and less interactive students more than it did comparatively better-performing students.
[…]
The universities involved in the study are constituent universities of the same university system with similar standards. The instructors of the two institutions coordinated on content taught, learning activities, assessments, and grading policies and standards. Students’ scores in each assignment were collected in Canvas and shared among instructors. The grade sheets can be exported and made available, with students’ personal information redacted. Statistical calculations were performed using Excel and Python packages. The findings of this study, concluding that effective note-taking helps in academic excellence, are consistent with those of other studies examining the role of effective note-taking in learning.
@@ V Statistical Tests and Limitation of the Study
We performed a $\chi^{2}$ goodness-of-fit test to determine whether the distributions of exam-only scores and aggregated scores were statistically similar. The scores were grouped into custom ranges: below 60, 60–70, 70–80, 80–90, and 90–100, and the $\chi^{2}$ values were evaluated using these categories as
$\chi^{2}=\sum_{i=1}^{n}\frac{(E_{i}-A_{i})^{2}}{E_{i}}\,,$ (1)
where $E_{i}$ and $A_{i}$ are frequencies for exam only and aggregated scores in $i^{\rm th}$ interval. The $\chi^{2}$ values for Phys-I and Phys-II are 23.9 and 24.5 respectively, indicating that the frequency distributions differ significantly. The $p$-values in each case is less than 0.05 suggesting that the difference in these two distributions in each course is statistically significant and they did not occur due to random chance. At a significance level of $\alpha=0.05$ (or 95% confidence interval), the statistical power is about 70%, meaning there is a 70% chance of correctly detecting a true effect. In statistical hypothesis testing, this is known as a Type II error, also called a false negative [52].
Figure 4: Plot showing the required sample size for 80% statistical power for as a function of effect size in terms of Cramér’s $V$-values. For the $V$-values of the students’ grade sheets presented in this study, the required sample size for Phys-I to get 80% statistical power is 39 and that for Phys-II is 45, which are slightly larger than the sizes in this study.
@@
where $n$ and $k$ are total number of observations and number of categories respectively. The Cramér’s $V$ values in Phys-I data is 0.45 while that for Phys-II is 0.42. To reliably reject a false null hypothesis, a statistical power of 80% or higher is commonly recommended at a significance level of $\alpha=0.05$ [53]. To increase reliability, a larger sample size would help.
The study was conducted at only two institutionseach offering a limited number of introductory physics course sections. Expanding the study to include multiple institutions and a larger student population would enhance the reliability of the results. Reliability could be improved by conducting the study with half of the students in a class using note-taking strategies, while the other half do not, in which students could self-select their group or be randomly assigned by the instructor. Instead, one can implement these two strategies in different chapters/modules for a chosen course/section. Alternatively, one can investigate the role of note-taking by implementing the above-mentioned strategies to assess the same group of students across different midterm exams.
@@ VI Conclusion
The impact of effective note-taking on learning in undergraduate-level algebra-based introductory physics courses at minority-serving universities was studied. We revisited the importance of a distraction-free and student-centered active learning environment to enhance the retention of learning materials in long-term memory, which students can retrieve effortlessly as needed and apply the received knowledge, concepts, and gained skills in various contexts now and in the future. A few commonly used note-taking methods were discussed. As an example of the mind mapping method of note-taking, we have presented a mind map to discuss techniques that students may follow to increase their attentiveness in the classroom, making physics learning fruitful. We discussed various approaches we have implemented in teaching and assessing algebra-based undergraduate-level introductory physics courses, as well as the impact of our efforts to make physics learning fun, active, student-centered, interactive, and objective-oriented.
We found that the majority of students were motivated to prepare better notes and actively engage in the classroom, as they could use their handwritten notes in timed assignments and exams. The final letter grades that students received at the end of the semester were recorded as a quantitative measurement of learning effectiveness. The letter grade significantly improved when students were allowed to prepare and use the notes they prepared during exams and timed assignments. Not only did the percentage of students getting higher grades increase with note-taking implementation, but the percentage of students failing the course also decreased sharply. The study revealed that implementing a note-taking learning approach comparatively enhances the grades of low—and average-performing students more than those of high-performing students.
The findings recommend that educators consider effective note-taking as an integral part of learning activities. In the current age of digital distraction, it is more important than ever. Asking students to make their mind map connecting different concepts in all the modules (chapters) covered in a course as note-taking homework, which we encouraged students to prepare, is excellent homework any instructor can assign. The mind mapping diagram presented in the manuscript serves as a strong example.

### v3

@@ III.2 Study Design
FSU follows the FN grade policy, which means that if students do not attend the first week of class, they will automatically drop out after the instructor notifies the registration office. There was no mid-semester withdrawal. Two students in the sections implementing no notetaking and one student from sections where notetaking was implemented did not take the midterm exam. Students dropped due to FN grades, and those who did not take the exam are not included in the analysis.
@@
Keeping the other factors the same, teaching pedagogy differing with the preparation and use of notetaking in two sets of sections was implemented. Instructor taught effective notetaking methods, asked students to prepare handwritten notes, and let them use the notes during classwork, timed quizzes, and exams. Study shows that hand-written notes increase comprehension, create a personal connection to the learning materials, improve memory and recall of the facts, and activate brain memory areas more robustly than preparing notes using digital devices, showing the pen mightier than the keyboard, although taking notes on electronic devices is faster, and the note taker can write to a larger extent of word-for-word what the teacher is saying in a class [52]. Students using electronic devices such as iPads, tablets, and laptops to take notes are usually prone to copying, pasting, and collecting information without a deep thought about what they are writing. In contrast, hand-written note-takers usually mentally process the information further before they write it down; while they write, they get their fingers’ extra muscle memory. Therefore, this research is designed as a comparative study of the effectiveness of handwritten notes in enhancing attention and effectiveness in physics learning. In order to maintain consistency, the quizzes and exams administered for both groups were problem-solving based, which demand critical thinking, creativity, and decision-making. Keeping all other conditions and class structure unaltered, the methodology was split into two main categories.
@@
      This policy was implemented in the first introductory physics lecture course, where the notetaking strategy was not discussed, nor were students allowed to use their notes in the exams. However, a list of all required mathematical expressions covered in the chapters, aka the formula sheet, was provided during timed assignments and exams. Scores from a terminal (third) exam are analyzed.
@@
    Handwritten notetaking is encouraged and permitted for use during exams:
      The notetaking policy was implemented in the first introductory physics course. At the beginning of the semester, a few samples of the notetaking skeleton were provided, and instructors consistently observed students’ notes during class to assess their effectiveness and guided them as needed. Only the handwritten notes prepared by students were permitted during exams. The notes students used in the timed quizzes and exams were monitored. The scores students received in a terminal (third) exam are analyzed.
In the sections where the notetaking strategy was implemented, a few samples of skeleton notes of each kind discussed above, along with sample notes, were also provided in the very first week of the semester. In addition, students received timely feedback on their notes, and those who struggled with notetaking and other academic things were encouraged to use office hours for additional support. The instructor had at least eight regularly scheduled office hours per week to help students with various aspects, including course content, academic and career plans, and other academic support. The instructor asked students to make mind maps connecting different concepts in all the chapters covered as notetaking homework, which the instructor found very helpful for students’ overall understanding of the course.
@@ IV Results and discussion
The grade that each student receives is the grade that students earn based on their academic performance, not measured against the performance of other students. The letter grades are assigned according to the University’s catalog as tabulated in Table 1. At FSU, using the traditional five-grade system, a letter grade of ”A” (GPA of 4.0 out of 4.0) is assigned if a student receives 90% of possible credits or more than that. Letter grades ”B” (GPA of 3.0 out of 4.0), ”C” (GPA of 2.0 out of 4.0), and ”D” (GPA of 1.0 out of 4.0) are assigned if a student’s grade point lies in the range of 80 to 89.9%, 70 to 79.9%, and 70 to 79.9%, respectively. A student fails the course if the student receives less than 60% of the total possible points.
@@
Figure 3: Bar chart showing letter grade percentage in sections with no notetaking implemented and with notetaking implemented. The grade presented in this bar chart is not the final grade student received, instead it is the grade students received in a mid-term exam. There were a total of 69 students involved in the teaching approach with no notetaking and total of 66 students in the notetaking implemented group.
@@
Table 2: Statistical measures of students’ performance in two different approaches to teaching introductory physics course in a terminal exam. The highest, lowest, and average scores were presented on the scale of 0–100.
Statistical measures Without notetaking With notetaking
Sample size 69 66
Highest score 100 100
Lowest score 31 55
Average score 75.3 85.4
Standard deviation 14.4 11.0
Figure 3 shows the percentage of letter grades that students received in sections with no notetaking and sections with notetaking were implemented. The sample size for these two groups in the study is almost the same, one being 69 and the other being 66. The data shows that when students were allowed to take notes, the percentage of students receiving the highest letter grade (i.e., ”A”) increased significantly, the second-highest letter grade (i.e., ”B”) increased slightly, while the lower letter grades decreased significantly. The percentage of students getting the letter grade ”A” is more than triple with the implementation of notetaking (please see and compare the pair of bars for letter grade ”A”). More than 40% of students mastering the content and getting ”A” grades indicate that the notetaking strategy works well in learning physics. Students receiving letter grades ”C” and ”D” and failing grade percentages dropped significantly when notetaking was implemented. Without notetaking, the letter grade peaked at ”C” while the letter grade peaked at ”A” and tailed to a lower grade when notetaking is implemented. The percentage of students who failed the course decreased from 14.3% to only 4.6%.
The statistical measures of student performance while the two different notetaking policies were implemented in teaching are presented in Table 2. Two out of 69 and one out of 66 students received perfect scores in the no-notetaking and with-notetaking groups, respectively. The average score in the class improved significantly ($\sim 10\%$) in the sections that used handwritten notes, although there was a lowest score outlier of 31. The standard deviation, which measures the sparseness of scores about the average score, decreased in the sections where the notetaking policy was implemented, indicating overall performance improvement in those sections. The findings of this study, concluding that effective notetaking helps in academic excellence, are consistent with those of other studies examining the role of effective notetaking in learning. This study also affirms the concept of cognetive load theoy that instructional methods can be optimized to improve understanding and retain knowledge longer [55].
@@ V Statistical Tests and Limitation of the Study
At a significance level of $\alpha=0.05$ (or 95% confidence interval), the statistical power is about 70%, meaning there is a 70% chance of correctly detecting a true effect. In statistical hypothesis testing, this is known as a Type II error, also called a false negative [53]. To reliably reject a false null hypothesis, a statistical power of 80% or higher is commonly recommended at a significance level of $\alpha=0.05$ [54].
Figure 4: Plot showing the required sample size for 80% statistical power for as a function of effect size in terms of Cramér’s $V$-values.
@@
where $\chi^{2}$, $n$ and $k$ are chi-square statistic, total number of observations and number of categories respectively. As shown in Fig. 4, the Cramér’s $V$ values for the sample size of 69, i.e, without note taking, is 0.335, while that for the sample size of 66, i.e., with note taking, is 0.345. The $\chi^{2}$ values of students’ scores in the group without notetaking is 8.90 with 4 degrees of freedom. The same for notetaking implemented group is 35.52 with 4 degrees of freedom. The Cramér’s $V$ value for the dataset for the notetaking not implemented group is 0.180, showing a weak association and suggesting that students’ score distribution differs only slightly from the uniform distribution. On the other hand, Cramér’s $V$ value for the notetaking implemented group is 0.367, indicating a moderate association strength and indicating scores concentrated in the higher bins 80-90 and 90-100, i.e., higher letter grades.
The study was conducted at only one institution offering a limited number of introductory physics course sections. Expanding the study to include multiple institutions and a larger student population would enhance the reliability of the results. Reliability could be improved by conducting the study with half of the students in a class using notetaking strategies, while the other half do not, in which students could self-select their group or be randomly assigned by the instructor. Alternatively, one can investigate the role of notetaking by implementing the above-mentioned strategies to assess the same group of students across different midterm exams. Furthermore, FSU is a historically black university in which the majority of students enrolled are from historically underserved groups. Carrying out the same or similar study in a diverse population would increase reliability of the research findings.
We acknowledge the fact that “grades” are considered the outcome. In many contexts, grades are short-term measures, while conceptual retention is a long-term measure. Conceptual retention over the years is out of the scope of this study. One may want to study the effect of notetaking measures in various other settings, such as larger universities, community colleges, and non-STEM courses, and compare the outcomes with those presented in this study.
@@ VI Conclusion
This study investigates the impact of effective notetaking on learning outcomes in undergraduate-level introductory physics courses at a minority-serving university. We highlighted the importance of a distraction-free, student-centered active learning environment, which enhances the retention of learning materials in long-term memory. This approach enables students to retrieve and apply the knowledge, concepts, and skills that they have acquired in various contexts both now and in the future. A few commonly used notetaking methods were discussed. As an example of the mind mapping method of notetaking, we have presented a mind map to discuss techniques that students may follow to increase their attentiveness in the classroom, making physics learning fruitful. We discussed various approaches we have implemented in teaching and assessing algebra-based undergraduate-level introductory physics courses, as well as the impact of our efforts to make physics learning fun, active, student-centered, interactive, and objective-oriented.
We found that the majority of students were motivated to prepare better notes and actively engage in the classroom, as they could use their handwritten notes in timed assignments and exams. The letter grades that students received for a mid-term exam were recorded as a quantitative measurement of learning effectiveness. The letter grade significantly improved when students were allowed to prepare and use the notes they prepared during exams and timed assignments. Not only did the percentage of students getting higher grades increase with notetaking implementation, but the percentage of students failing the course also decreased sharply.
The findings recommend that educators consider effective notetaking as an integral part of learning activities.In today’s digital age, minimizing distraction is crucial. Asking students to make their mind map connecting different concepts in all the modules (chapters) covered in a course as notetaking homework, which we encouraged students to prepare, is an excellent homework any instructor can assign.
Institutional Review
This study was reviewed and approved by the Human Rights in Research Committee (HRRC) at Fayetteville State University with the assigned IRB #2026-20 under ”Exempt Status”.
Acknowledgements.
The author acknowledges the communications with Drs. Tikaram Neupane and Uma Poudyal.